The number of emails that I get in one hour has such density function:
x 2 1 4 5 3 6 and more 0
f(x) 0.17 0.23 0.08 0.04 0.14 0.02 ?
a) find the probability, that there will be no letter for the next hour;
b) find cumulative distributional function.
1
Expert's answer
2020-04-10T18:26:51-0400

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c:c}
 x & 0 & 1 & 2 & 3 & 4 & 5 & 6\text{ and more} \\ \hline
 f(x) & a & 0.23 & 0.17& 0.14 & 0.08 & 0.04 & 0.02 \\
 
\end{array}"

a) "\sum \limits _{x} P(X=x)=1 \ \ \Rightarrow \ \ a+0.23+0.17+0.14+0.08+0.04+0.02=1 \ \ \Rightarrow \ \ a=0.32"

Answer: the probability, that there will be no letter for the next hour, is equal to 0.32


b)

The cumulative distribution function is given by "F(x)=P(X\leq x)"

"P(X\leq x)=\sum \limits _{y\leq x}P(X=y)= \sum \limits _{y\leq x-1}P(X=y)+P(X=x)=P(X\leq x-1)+P(X=x)"

"F(x)=F(x-1)+P(X=x) \ \ (\text{for }x>0)"

Now we can calculate "F(x):"

"F(0)=0.32"

"F(1)=0.32+0.23=0.55"

"F(2)=0.55+0.17=0.72"

"F(3)=0.72+0.14=0.86"

"F(4)=0.86+0.08=0.94"

"F(5)=0.94+0.04=0.98"

"F(6 \text{ and more}) =0.98+0.02=1"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c:c}
 x & 0 & 1 & 2 & 3 & 4 & 5 & 6\text{ and more} \\ \hline
 F(x) & 0.32 & 0.55 & 0.72& 0.86 & 0.94 & 0.98 & 1 \\
 
\end{array}"